40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Engineering group Lab elog  Not logged in ELOG logo
Message ID: 253     Entry time: Fri Sep 3 14:25:00 2021
Author: Stephen 
Type: Progress 
Category: RTS for COC 
Subject: Anneal of SN0932 at 300C for 10 hours 

SN0932 was annealed between 02 and 03 September 2021 - used the large furnace in Downs 221, atmosphere was air, used glass petri dishes to protect optic (from Gabriele and CRiMe).

Controller parameters:

Ramp up rate = 100 °C per hour
Hold temp = 300 °C
Hold time = 10 hours
Ramp down rate = 100 °C per hour

Witness RTD observed a ~22 °C overshoot and ~14 °C offset from controller RTD. See attached image (.xlsx file is source document).

Optic is now with Liyuan in RTS for recharacterization after anneal.

 

Attachment 1: anneal_sn0932_300C_20210903.png  25 kB  Uploaded Fri Sep 3 16:35:21 2021  | Hide | Hide all
anneal_sn0932_300C_20210903.png
Attachment 2: anneal_sn0932_300C_20210903.xlsx  85 kB  Uploaded Fri Sep 3 16:40:08 2021
ELOG V3.1.3-